# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 6 Application: Over-parameterized Matrix Sensing
In this section, we apply our general framework the over-parameterized matrix sensing. In this problem, the aim is to recover an $n\times n$ and rank-$r$ PSD matrix ${\mathbf{M}^{\star}}$ from a set of linear measurements $\{(y_{i},\mathbf{A}_{i})\}_{i=1}^{N}$, where $y_{i}=\left\langle\mathbf{A}_{i},{\mathbf{M}^{\star}}\right\rangle$ for $i=1,\dots,N$. To this end, a common approach is to employ a factorized model ${\mathbf{M}}=\mathbf{X}\mathbf{X}^{\top}$, where $\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}$ and $r^{\prime}\geq r$, and minimize the following mean-squared error
@@
When the search rank $r^{\prime}$ exceeds the true rank $r$, the problem is over-parameterized. For simplicity, we define the linear operator $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{N}$ and its adjoint $\mathcal{A}^{*}:\mathbb{R}^{N}\to\mathbb{R}^{n\times n}$ as $\mathcal{A}(\mathbf{M})=\frac{1}{\sqrt{N}}\begin{bmatrix}\left\langle\mathbf{A}_{1},\mathbf{M}\right\rangle&\cdots&\left\langle\mathbf{A}_{N},\mathbf{M}\right\rangle\end{bmatrix}^{\top}$ and $\mathcal{A}^{*}(\mathbf{y})=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}y_{i}\mathbf{A}_{i}$, respectively. A typical assumption on $\mathcal{A}$ is that it satisfies the so-called restricted isometry property (RIP), defined below.
@@
$(1-\delta)\left\lVert\mathbf{X}\right\rVert_{\mathrm{F}}^{2}\leq\left\lVert\mathcal{A}(\mathbf{X})\right\rVert^{2}\leq(1+\delta)\left\lVert\mathbf{X}\right\rVert_{\mathrm{F}}^{2},\quad\text{for all $\mathbf{X}\in\mathbb{R}^{n\times n}$ with $\mathrm{rank}(\mathbf{X})\leq r$}.$
@@
Intuitively, the RIP condition requires the operator $\mathcal{A}$ to be nearly isotropic over all low-rank matrices, a property that holds when $\mathcal{A}$ exhibits a certain type of randomness. For instance, it is well-known that for a Gaussian operator $\mathcal{A}$, where all the elements of the matrices $\mathbf{A}_{i}$ are independently drawn from a standard Normal distribution, $(\delta,r)$-RIP holds with a high probability, provided that the number of measurements satisfies $N=\Omega\left({dr}/{\delta^{2}}\right)$ [54].
To apply our proposed framework, we first need to define the implicit region $\mathcal{M}$. To this end, we denote the nonzero singular values of ${\mathbf{M}^{\star}}$ in descending order as $\sigma_{1}^{\star}\geq\cdots\geq\sigma_{r}^{\star}>0$, with $\kappa=\sigma_{1}^{\star}/\sigma_{r}^{\star}$ representing the condition number. The singular value decomposition (SVD) of ${\mathbf{M}^{\star}}$ is given by ${\mathbf{M}^{\star}}=\mathbf{V}^{\star}\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}$, where $\bm{\Sigma}^{\star}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the nonzero eigenvalues of ${\mathbf{M}^{\star}}$, and $\mathbf{V}^{\star}\in\mathcal{O}_{n\times r}$ is the matrix of corresponding eigenvectors. Moreover, let $\mathbf{V}^{\star\perp}\in\mathcal{O}_{n\times(n-r)}$ denote the orthogonal complement of $\mathbf{V}^{\star}$. We define the implicit region as follows:
@@
$\mathcal{M}$ =$=\left\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r,\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta,\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}\right\}.$ (78)
Intuitively, the implicit region above corresponds to the set of rank-$r$ matrices whose column spaces do not deviate significantly from that of the true solution, and whose norms are bounded from above. We also note that every true solution of the form $\mathbf{X}^{\star}\mathbf{X}^{\star\top}={\mathbf{M}^{\star}}$ can be equivalently characterized as $\mathcal{X}^{\star}=\{\mathbf{X}:\mathbf{X}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}^{\top},\mathbf{O}\in\mathcal{O}_{r^{\prime}\times r}\}$ [55, Lemma 20]. This implies that $\mathcal{X}^{\star}\subset\mathcal{M}$. Moreover, $\mathbf{X}$ is $\mathcal{M}$-SOSP if and only if $\mathbf{X}\in\mathcal{X}^{\star}$—that is, $\mathcal{X}_{\mathcal{M}}=\mathcal{X}^{\star}$—provided that the linear operator $\mathcal{A}$ satisfies $(\delta,5r)$-RIP with $\delta\leq\frac{1}{10}$ [49, Theorem 37]. Therefore, to approximately recover the true rank-$r$ solution from $\mathcal{X}^{\star}$, it suffices to recover an approximate $\mathcal{M}$-SOSP.
@@
Toward this goal, we show that the conditions of Theorem 8 are guaranteed to hold, leading to a nearly linear convergence of IPGD to an $\mathcal{M}$-SOSP. To this goal, first, we introduce an important implication of the RIP condition, which will be used extensively throughout this section.
Lemma 22 (Lemma 2.3. and Lemma C.2. in [5]).
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,r)$-RIP. Then for any matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$, we have $\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)(\mathbf{M})\right\rVert\leq\delta\left\lVert\mathbf{M}\right\rVert_{\star}$, where $\left\lVert\mathbf{M}\right\rVert_{\star}$ is the nuclear norm of $\mathbf{M}$. Moreover, if $\mathrm{rank}(\mathbf{M})\leq r$, we further have $\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)(\mathbf{M})\right\rVert\leq\sqrt{r}\delta\left\lVert\mathbf{M}\right\rVert$.
For simplicity, we denote $\bm{\Delta}={\mathbf{M}^{\star}}-\mathbf{X}\mathbf{X}^{\top}$.
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,2r)$-RIP with $\delta\leq\frac{1}{10\sqrt{r}}$. With the implicit region defined as (78), the function $f(\mathbf{X})$ is ${(L,\mathcal{M},\tau)}$-gradient-Lipschitz and ${(\rho,\mathcal{M},\tau)}$-Hessian-Lipschitz, with parameters $L=15\sigma_{1}^{\star}$, $\rho=15\sqrt{\sigma_{1}^{\star}}$, and $\tau=\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$.
@@
The first formula follows from [42, Lemma 7]. We omit the derivation of the second formula, as it can be obtained in a similar manner. To derive the gradient and Hessian Lipschitz constants, we use the following equivalent definitions for smooth functions [56, Theorem 9.19]:
@@
We first provide an upper bound for $L$. For any $\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)$ and $\left\lVert\mathbf{Z}\right\rVert_{\mathrm{F}}\leq 1$, Equation 79 implies
@@
where in the last inequality, we use triangle inequality and the fact that $\left\lVert\mathbf{Z}\right\rVert_{\mathrm{F}}\leq 1$ repeatedly. Note that $\mathbf{X}\in\mathcal{N}_{\mathcal{M}}(\tau)$ implies that $\mathrm{rank}\big{(}\mathbf{X}^{\sharp}\big{)}\leq r$, $\left\lVert\mathbf{X}^{\sharp}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$, and $\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\leq\tau$. Consider the following decomposition:
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sqrt{2r}\delta\big{\|}\bm{\Delta}^{\sharp}\big{\|}+\delta\big{\|}\bm{\Delta}^{\perp}\big{\|}_{\star}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}4\sqrt{2r}\sigma_{1}^{\star}\delta+\delta\cdot 6\sqrt{d\sigma_{1}^{\star}}\tau$
$\stackrel{{\scriptstyle(c)}}{{\leq}}10\sqrt{r}\sigma_{1}^{\star}\delta,$
@@
where we use the fact that $\delta\leq\frac{1}{10\sqrt{r}}$. This implies that the objective function is gradient Lipschitz with Lipschitz constant $L=15\sigma_{1}^{\star}$ within $\mathcal{N}_{\mathcal{M}}(\tau)$. In a similar fashion, we can show that the objective function is Hessian Lipschitz with Lipschitz constant $\rho=15\sqrt{\sigma_{1}^{\star}}$ within $\mathcal{N}_{\mathcal{M}}(\tau)$. We omit the details for brevity. ∎
Closure of $\mathcal{M}$ under GD updates
Next, we establish the closure of $\mathcal{M}$ under GD updates.
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,2r)$-RIP with $\delta\leq\frac{1}{10\sqrt{r}}$. For any $\mathbf{X}\in\mathcal{M}$, we have $\mathbf{X}^{+}:=\mathbf{X}-\eta\nabla f(\mathbf{X})\in\mathcal{M}$, provided that the step-size satisfies $\eta\leq\frac{1}{10\sigma_{1}^{\star}}$.
@@
Lemma 25 (Adapted from [57, Lemma 2.1.2 and Lemma 2.1.3]).
@@
$\left\lVert\mathcal{P}_{\mathbf{U}^{\perp}}\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert=\left\lVert\mathcal{P}_{\mathbf{U}}-\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert\leq\min_{\mathbf{R}\in\mathcal{O}^{r\times r}}\left\lVert\mathbf{U}\mathbf{R}-\mathbf{U}^{\star}\right\rVert\leq\sqrt{2}\left\lVert\mathcal{P}_{\mathbf{U}}-\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert=\sqrt{2}\left\lVert\mathcal{P}_{\mathbf{U}^{\perp}}\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert.$ (87)
Lemma 26 (Adapted from [58, Theorem 2.4]).
@@
$\mathbf{X}^{+}:=\mathbf{X}-\eta\nabla f(\mathbf{X})=\Big{(}\mathbf{I}-\eta\mathcal{A}^{*}\mathcal{A}\big{(}\mathbf{X}\mathbf{X}^{\top}-{\mathbf{M}^{\star}}\big{)}\Big{)}\mathbf{X}=(\mathbf{I}+\eta\mathcal{A}^{*}\mathcal{A}(\bm{\Delta}))\mathbf{X}.$ (90)
Based on the definition $\mathcal{M}$ in Equation 78, to show that $\mathbf{X}^{+}=(\mathbf{I}+\eta\mathcal{A}^{*}\mathcal{A}(\bm{\Delta}))\mathbf{X}\in\mathcal{M}$, we need to verify the following three conditions
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\left(1+\eta\left(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-\left\lVert\mathbf{X}\right\rVert^{2}\right)\right)\left\lVert\mathbf{X}\right\rVert$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left(1+\eta\left(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-4\sigma_{1}^{\star}\right)\right)\cdot 2\sqrt{\sigma_{1}^{\star}}$
$\leq 2\sqrt{\sigma_{1}^{\star}}.$
In $(a)$, we use triangle inequality and Lemma 22 along with the facts that $\mathrm{rank}\left(\bm{\Delta}\right)\leq 2r$ and $\left\lVert\bm{\Delta}\right\rVert\leq 4\sigma_{1}^{\star}$. In $(b)$, we use the fact that the function $f(x)=x-\eta x^{3}$ is increasing in the interval $[0,\frac{1}{\sqrt{3\eta}}]$, which implies $\left\lVert\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\mathbf{X}\right\rVert=\left\lVert\mathbf{X}\right\rVert-\eta\left\lVert\mathbf{X}\right\rVert^{3}$ since $\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}\leq\frac{1}{\sqrt{3\eta}}$. Lastly, in $(c)$, we use the fact that the function $g(x)=(1+\eta(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-x^{2}))x$ is also increasing in the range $[0,\frac{1}{3\sqrt{\eta}}]$.
@@
$\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\tilde{\mathbf{X}}^{+}}$ $=\mathcal{P}_{\mathbf{V}^{\star\perp}}\tilde{\mathbf{X}}^{+}(\tilde{\mathbf{X}}^{+})^{\dagger}$ (93)
$=\mathcal{P}_{\mathbf{V}^{\star\perp}}\left(\mathbf{I}+\eta\bm{\Delta}\right)\mathbf{X}(\tilde{\mathbf{X}}^{+})^{\dagger}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathcal{P}_{\mathbf{V}^{\star\perp}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\mathbf{X}(\tilde{\mathbf{X}}^{+})^{\dagger}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)(\mathbf{I}+\eta\bm{\Delta})^{-1}\mathcal{P}_{\tilde{\mathbf{X}}^{+}},$
where, in $(a)$, we use the fact that $\mathcal{P}_{\mathbf{V}^{\star\perp}}{\mathbf{M}^{\star}}=\bm{0}$. Moreover, in $(b)$, we use the fact that $\mathbf{X}(\tilde{\mathbf{X}}^{+})^{\dagger}=(\mathbf{I}+\eta\bm{\Delta})^{-1}\tilde{\mathbf{X}}^{+}(\tilde{\mathbf{X}}^{+})^{\dagger}=(\mathbf{I}+\eta\bm{\Delta})^{-1}\mathcal{P}_{\tilde{\mathbf{X}}^{+}}$. Therefore, we obtain

### v2

@@ 6 Application: Over-parameterized Matrix Sensing
In this section, we apply our general framework to the over-parameterized matrix sensing problem. As highlighted in Section 1.2, general-purpose saddle-avoiding methods such as PGD fail to converge to the true solution in this setting. In contrast, we prove the convergence of IPGD for this problem.
Recall that here, the aim is to recover an $n\times n$ and PSD matrix $\mathbf{\Theta}^{\star}$ of rank at most $r$ from a set of linear measurements $\{(y_{i},\mathbf{A}_{i})\}_{i=1}^{N}$, where $y_{i}=\left\langle\mathbf{A}_{i},\mathbf{\Theta}^{\star}\right\rangle$ for $i=1,\dots,N$. Therefore, the feasible region can be defined as $\mathcal{D}=\{\mathbf{\Theta}\in\mathbb{R}^{n\times n}:\mathbf{\Theta}\succeq 0,\mathrm{rank}(\mathbf{\Theta})\leq r\}$. To this end, a common approach is to employ a reparameterization map $\varphi(\mathbf{X})=\mathbf{X}\mathbf{X}^{\top}$, where $\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}$ and $r^{\prime}\geq r$, and minimize $f=L\circ\varphi$, where $L$ is the mean-squared error:
@@
For simplicity, we define the linear operator $\mathcal{A}:\mathbb{R}^{n\times n}\to\mathbb{R}^{N}$ and its adjoint $\mathcal{A}^{*}:\mathbb{R}^{N}\to\mathbb{R}^{n\times n}$ as $\mathcal{A}(\mathbf{\Theta})=\frac{1}{\sqrt{N}}\left[\left\langle\mathbf{A}_{1},\mathbf{\Theta}\right\rangle\ \cdots\ \left\langle\mathbf{A}_{N},\mathbf{\Theta}\right\rangle\right]^{\top}$ and $\mathcal{A}^{*}(\mathbf{y})=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}y_{i}\mathbf{A}_{i}$, respectively. A typical assumption on $\mathcal{A}$ is that it satisfies the so-called restricted isometry property (RIP), defined below.
@@
$(1-\delta)\left\lVert\mathbf{\Theta}\right\rVert_{\mathrm{F}}^{2}\leq\left\lVert\mathcal{A}(\mathbf{\Theta})\right\rVert^{2}\leq(1+\delta)\left\lVert\mathbf{\Theta}\right\rVert_{\mathrm{F}}^{2},\quad\text{for all $\mathbf{\Theta}\in\mathbb{R}^{n\times n}$ with $\mathrm{rank}(\mathbf{\Theta})\leq r$}.$
@@
Intuitively, the RIP condition requires the operator $\mathcal{A}$ to be nearly isotropic over all low-rank matrices, a property that holds when $\mathcal{A}$ exhibits a certain type of randomness. For instance, it is well-known that for a Gaussian operator $\mathcal{A}$, where all the elements of the matrices $\mathbf{A}_{i}$ are independently drawn from a standard normal distribution, $(\delta,r)$-RIP holds with a high probability, provided that the number of measurements satisfies $N=\Omega\left({dr}/{\delta^{2}}\right)$ [candes2011tight].
When $r$ is unknown, it is common to over-parameterize the model by choosing a search rank $r^{\prime}\leq n$ that exceeds the true rank $r$. Indeed, one common choice is $r^{\prime}=n$ (referred to as fully over-parameterized model). Under this assumption, we have $\mathcal{M}=\varphi^{-1}(\mathcal{D})=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r\}$. However, verifying the conditions of Theorem 3 or Theorem 4 using this implicit region is challenging for two key reasons. First, because of the quartic structure of $f$ and the unboundedness of $\mathcal{M}$, neither gradient- nor Hessian-Lipschitz condition holds on $\mathcal{M}$. Second, within this implicit region, it is difficult to establish a tight bound on the deviation rate.
We show that both of these challenges can be circumvented with a careful refinement of $\mathcal{M}$. Let the eigen-decomposition of $\mathbf{\Theta}^{\star}$ be given by $\mathbf{\Theta}^{\star}=\mathbf{V}^{\star}\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}$, where $\bm{\Sigma}^{\star}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the nonzero eigenvalues of $\mathbf{\Theta}^{\star}$, and $\mathbf{V}^{\star}\in\mathcal{O}_{n\times r}$ is the matrix of corresponding eigenvectors. We rely on the key result that all SOSPs within $\mathcal{M}$ coincide with $\mathcal{X}^{\star}=\{\mathbf{X}:\mathbf{X}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}^{\top},\ \mathbf{O}\in\mathcal{O}_{r^{\prime}\times r}\}$—that is, $\mathcal{X}_{\mathcal{M}}=\mathcal{X}^{\star}$—provided that the linear operator $\mathcal{A}$ satisfies the $(\delta,5r)$-RIP with $\delta\leq\tfrac{1}{10}$ [chi2019nonconvex, Theorem 37]. Since $\mathbf{X}^{\star}\mathbf{X}^{\star\top}={\mathbf{M}^{\star}}$ if and only if $\mathbf{X}^{\star}\in\mathcal{X}^{\star}$ [ma2023can, Lemma 20], to approximately recover the true rank-$r$ solution, it suffices to recover an approximate $\mathcal{M}$-SOSP. Leveraging this characterization of $\mathcal{M}$-SOSPs, we consider a restricted version of $\mathcal{M}$ that addresses both of the aforementioned challenges while still containing all the $\mathcal{M}$-SOSPs:
@@
$\mathcal{M}_{\text{RIP}}$ $=\left\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r,\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta,\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}\right\}.$ (79)
In the above equation, $\kappa=\sigma_{1}^{\star}/\sigma_{r}^{\star}$ represents the condition number of $\mathbf{\Theta}^{\star}$, where $\sigma_{1}^{\star}$ and $\sigma_{r}^{\star}$ are the largest and smallest nonzero eigenvalues of $\mathbf{\Theta}^{\star}$, respectively. Moreover, $\mathbf{V}^{\star\perp}\in\mathcal{O}_{n\times(n-r)}$ denotes the orthogonal complement of $\mathbf{V}^{\star}$. It is easy to verify that $\mathcal{X}^{\star}\subset\mathcal{M}_{\text{RIP}}$, implying that the SOSPs within $\mathcal{M}$ and $\mathcal{M}_{\text{RIP}}$ coincide. Intuitively, $\mathcal{M}_{\text{RIP}}$ corresponds to the set of rank-$r$ matrices whose column spaces do not deviate significantly from that of the true solution, and whose norms are bounded from above. As will be shown later, the former property enables us to establish sharp bounds on the deviation rate of GD, while the latter ensures that both the gradient- and Hessian-Lipschitz conditions hold on $\mathcal{M}_{\text{RIP}}$. Moreover, note that the zero matrix $\bm{0}_{n\times r^{\prime}}$ belongs to $\mathcal{M}_{\text{RIP}}$, and hence, can be used as a trivial initial point for the algorithm.
@@
Next, we show that the conditions of Theorem 4 are guaranteed to hold, leading to a nearly linear convergence of IPGD to an $\mathcal{M}_{\text{RIP}}$-SOSP. To this goal, first, we introduce an important implication of the RIP condition, which will be used extensively throughout this section.
Lemma 12 (Lemma 2.3. and Lemma C.2. in [li2018algorithmic]).
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,r)$-RIP. Then for any matrix $\mathbf{\Theta}\in\mathbb{R}^{n\times n}$, we have $\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)(\mathbf{\Theta})\right\rVert\leq\delta\left\lVert\mathbf{\Theta}\right\rVert_{\star}$, where $\left\lVert\mathbf{\Theta}\right\rVert_{\star}$ is the nuclear norm of $\mathbf{\Theta}$. Moreover, if $\mathrm{rank}(\mathbf{\Theta})\leq r$, we further have $\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)(\mathbf{\Theta})\right\rVert\leq\sqrt{r}\delta\left\lVert\mathbf{\Theta}\right\rVert$.
For simplicity, we denote $\bm{\Delta}=\mathbf{\Theta}^{\star}-\mathbf{X}\mathbf{X}^{\top}$.
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,2r)$-RIP with $\delta\leq\frac{1}{10\sqrt{r}}$. With the implicit region defined in (79), the function $f(\mathbf{X})$ is $(L,\mathcal{M}_{\text{RIP}},\tau)$}-gradient-Lipschitz and $(\rho,\mathcal{M}_{\text{RIP}},\tau)$}-Hessian-Lipschitz, with parameters $L=15\sigma_{1}^{\star}$, $\rho=15\sqrt{\sigma_{1}^{\star}}$, and $\tau=\frac{1}{10}\sqrt{\frac{\sigma_{1}^{\star}}{d}}$.
@@
The first formula follows from [ge2017no, Lemma 7]. We omit the derivation of the second formula, as it can be obtained in a similar manner. To derive the gradient and Hessian Lipschitz constants, we use the following equivalent definitions for smooth functions [rudin1964principles, Theorem 9.19]:
@@
We first provide an upper bound for $L$. For any $\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)$ and $\left\lVert\mathbf{Z}\right\rVert_{\mathrm{F}}\leq 1$, Equation 80 implies
@@
where in the last inequality, we use triangle inequality and the fact that $\left\lVert\mathbf{Z}\right\rVert_{\mathrm{F}}\leq 1$ repeatedly. Note that $\mathbf{X}\in\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)$ implies that $\mathrm{rank}\big(\mathbf{X}^{\sharp}\big)\leq r$, $\left\lVert\mathbf{X}^{\sharp}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}$, and $\left\lVert\mathbf{X}^{\perp}\right\rVert_{\mathrm{F}}\leq\tau$. Consider the following decomposition:
@@
[…]
@@
where we use the fact that $\delta\leq\frac{1}{10\sqrt{r}}$. This implies that the objective function is gradient Lipschitz with Lipschitz constant $L=15\sigma_{1}^{\star}$ within $\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)$. In a similar fashion, we can show that the objective function is Hessian Lipschitz with Lipschitz constant $\rho=15\sqrt{\sigma_{1}^{\star}}$ within $\mathcal{N}_{\mathcal{M}_{\text{RIP}}}(\tau)$. We omit the details for brevity.
Closure of $\mathcal{M}_{\text{RIP}}$ under GD updates
Next, we prove the closure of $\mathcal{M}_{\text{RIP}}$ under GD updates.
@@
Suppose that $\mathcal{A}$ satisfies $(\delta,2r)$-RIP with $\delta\leq\frac{1}{10\sqrt{r}}$. For any $\mathbf{X}\in\mathcal{M}_{\text{RIP}}$, we have $\mathbf{X}^{+}:=\mathbf{X}-\eta\nabla f(\mathbf{X})\in\mathcal{M}_{\text{RIP}}$, provided that the step-size satisfies $\eta\leq\frac{1}{10\sigma_{1}^{\star}}$.
@@
Lemma 13 (Adapted from [chen2021spectral, Lemma 2.1.2 and 2.1.3]).
@@
$\left\lVert\mathcal{P}_{\mathbf{U}^{\perp}}\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert=\left\lVert\mathcal{P}_{\mathbf{U}}-\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert\leq\min_{\mathbf{R}\in\mathcal{O}_{r\times r}}\left\lVert\mathbf{U}\mathbf{R}-\mathbf{U}^{\star}\right\rVert\leq\sqrt{2}\left\lVert\mathcal{P}_{\mathbf{U}}-\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert=\sqrt{2}\left\lVert\mathcal{P}_{\mathbf{U}^{\perp}}\mathcal{P}_{\mathbf{U}^{\star}}\right\rVert.$ (88)
Lemma 14 (Adapted from [chen2016perturbation, Theorem 2.4]).
@@
$\mathbf{X}^{+}:=\mathbf{X}-\eta\nabla f(\mathbf{X})=\Big(\mathbf{I}-\eta\mathcal{A}^{*}\mathcal{A}\big(\mathbf{X}\mathbf{X}^{\top}-\mathbf{\Theta}^{\star}\big)\Big)\mathbf{X}=(\mathbf{I}+\eta\mathcal{A}^{*}\mathcal{A}(\bm{\Delta}))\mathbf{X}.$ (91)
Based on the definition $\mathcal{M}_{\text{RIP}}$ in Equation 79, to show that $\mathbf{X}^{+}=(\mathbf{I}+\eta\mathcal{A}^{*}\mathcal{A}(\bm{\Delta}))\mathbf{X}\in\mathcal{M}_{\text{RIP}}$, we need to verify the following three conditions
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\left(1+\eta\left(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-\left\lVert\mathbf{X}\right\rVert^{2}\right)\right)\left\lVert\mathbf{X}\right\rVert$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left(1+\eta\left(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-4\sigma_{1}^{\star}\right)\right)\cdot 2\sqrt{\sigma_{1}^{\star}}\leq 2\sqrt{\sigma_{1}^{\star}}.$
In $(a)$, we use triangle inequality and Lemma 12 along with the facts that $\mathrm{rank}\left(\bm{\Delta}\right)\leq 2r$ and $\left\lVert\bm{\Delta}\right\rVert\leq 4\sigma_{1}^{\star}$. In $(b)$, we use the fact that the function $f(x)=x-\eta x^{3}$ is increasing in the interval $0,\frac{1}{\sqrt{3\eta}}]$, which implies $\left\lVert\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\mathbf{X}\right\rVert=\left\lVert\mathbf{X}\right\rVert-\eta\left\lVert\mathbf{X}\right\rVert^{3}$ since $\left\lVert\mathbf{X}\right\rVert\leq 2\sqrt{\sigma_{1}^{\star}}\leq\frac{1}{\sqrt{3\eta}}$. Lastly, in $(c)$, we use the fact that the function $g(x)=(1+\eta(\sigma_{1}^{\star}+4\sqrt{2r}\sigma_{1}^{\star}\delta-x^{2}))x$ is also increasing in the range $0,\frac{1}{3\sqrt{\eta}}]$.